### v1

# Entanglement witnesses and separability criteria based on generalized equiangular tight frames

Abstract
We use operators from generalized equiangular measurements to construct positive maps. Their positivity follows from the inequality for indices of coincidence corresponding to few equiangular tight frames. These maps give rise to entanglement witnesses, which include as special cases many important classes considered in the literature. Additionally, we introduce separability criteria based on the correlation matrix and analyze them for various types of measurements.
1 Measurements from generalized equiangular tight frames
A set $\{P_{k}:\,k=1,\ldots,M\}$ of rank-1 projectors $P_{k}$ acting on the Hilbert space $\mathcal{H}\simeq\mathbb{C}^{d}$ is an equiangular tight frame if $\mathrm{Tr}(P_{k}P_{\ell})=c\mathrm{Tr}(P_{k})\mathrm{Tr}(P_{\ell})$ for all $k\neq\ell$ and $\sum_{k=1}^{M}P_{k}=\gamma\mathbb{I}_{d}$ with $\gamma>0$ Strohmer . If $P_{k}$ are no longer rank-1, then they form a generalized equiangular tight frame GEAM . Now, take collections of $N$ generalized equiangular tight frames $\mathcal{P}_{\alpha}=\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha}\}$, $\alpha=1,\ldots,N$, whose elements sum up to $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$, where $\gamma_{\alpha}$ is a probability distribution. Additionally, assume that they are complementary PetzRuppert , which means that $\mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell})$, $\alpha\neq\beta$. Under these conditions, one defines the corresponding positive, operator-valued measure (POVM).
Definition 1.
(GEAM ) A collection of $N$ generalized equiangular lines $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}$ is a generalized equiangular measurement provided that
$\begin{split}\mathrm{Tr}(P_{\alpha,k})&=a_{\alpha},\\ \mathrm{Tr}(P_{\alpha,k}^{2})&=b_{\alpha}\mathrm{Tr}(P_{\alpha,k})^{2},\\ \mathrm{Tr}(P_{\alpha,k}P_{\alpha,\ell})&=c_{\alpha}\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\alpha,\ell}),\qquad k\neq\ell,\\ \mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})&=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell}),\qquad\alpha\neq\beta,\end{split}$ (1)
where
$a_{\alpha}=\frac{d\gamma_{\alpha}}{M_{\alpha}},\qquad c_{\alpha}=\frac{M_{\alpha}-db_{\alpha}}{d(M_{\alpha}-1)},\qquad f=\frac{1}{d},$ (2)
$\frac{1}{d}<b_{\alpha}\leq\frac{1}{d}\min\{d,M_{\alpha}\}.$ (3)
Generalized equiangular measurements can be constructed from Hermitian orthonormal bases $\{G_{0}=\mathbb{I}_{d}/\sqrt{d},G_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ with traceless $G_{\alpha,k}$. Namely, one has GEAM
$P_{\alpha,k}=\frac{a_{\alpha}}{d}\mathbb{I}_{d}+\tau_{\alpha}H_{\alpha,k},$ (4)
where the parameter
$\tau_{\alpha}^{2}=\frac{a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})}{M_{\alpha}(\sqrt{M_{\alpha}}+1)^{2}}$ (5)
is chosen this way to guarantee $P_{\alpha,k}\geq 0$. The traceless operators $H_{\alpha,k}$ are related to $G_{\alpha,k}$ in the following way,
$H_{\alpha,k}=\left\{\begin{aligned} &G_{\alpha}-\sqrt{M_{\alpha}}(1+\sqrt{M_{\alpha}})G_{\alpha,k},\quad k=1,\ldots,M_{\alpha}-1,\\ &(1+\sqrt{M_{\alpha}})G_{\alpha},\qquad k=M_{\alpha},\end{aligned}\right.$ (6)
where $G_{\alpha}=\sum_{k=1}^{M_{\alpha}-1}G_{\alpha,k}$.
There exists an important class of generalized equiangular measurements that are conical 2-designs Graydon ; Graydon2 – that is, they satisfy
$\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}\otimes P_{\alpha,k}=\kappa_{+}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\kappa_{-}\mathbb{F}_{d}$ (7)
for real parameters $\kappa_{+}\geq\kappa_{-}>0$, where $\mathbb{F}_{d}=\sum_{m,n=1}^{d}|m\rangle\langle n|\otimes|n\rangle\langle m|$ is the flip operator. It has been shown in ref. GEAM that a linear dependence between purity $\mathrm{Tr}(\rho^{2})$ of a state $\rho$ and the corresponding index of coincidence Rastegin5
$\mathcal{C}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad p_{\alpha,k}=\mathrm{Tr}(P_{\alpha,k}\rho),$ (8)
exists if and only if $P_{\alpha,k}$ form a conical 2-design.
Proposition 1.
The generalized equiangular measurement whose parameters satisfy $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, where
$0<S\leq\min\left\{\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\right\},$ (9)
is a conical 2-design with real parameters
$\kappa_{+}=\mu-\frac{S}{d},\qquad\kappa_{-}=S,\qquad\mu=\frac{1}{d}\sum_{\alpha=1}^{N}a_{\alpha}\gamma_{\alpha}.$ (10)
Observe that each generalized equiangular tight frame has its own index of coincidence (after rescaling)
$\mathcal{C}_{\alpha}=\frac{1}{\gamma_{\alpha}^{2}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad\frac{1}{\gamma_{\alpha}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=1.$ (11)
Therefore, it is possible to derive an inequality for sums of ssquared probabilities, similar to that for few mutually unbiased bases Wu .
Proposition 2.
If $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a conical 2-design, then
$\mathcal{C}_{L}=\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq S\left(\mathrm{Tr}\rho^{2}-\frac{1}{d}\right)+\mu_{L},$ (12)
where $\mu_{L}=(1/d)\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}$ and $1\leq L\leq N$. The equality is reached for $L=N$.
Proof.
In this proof, we follow the method introduced in ref. Rastegin2 . Any quantum state $\rho$ can be represented as
$\rho=\frac{1}{d}\mathbb{I}_{d}+\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}H_{\alpha,k}$ (13)
with real-valued parameters $r_{\alpha,k}$. Using the properties of the Hermitian operators $H_{\alpha,k}$,
$\begin{split}\mathrm{Tr}(H_{\alpha,k}^{2})&=(M_{\alpha}-1)(\sqrt{M_{\alpha}}+1)^{2},\\ \mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})&=-(\sqrt{M_{\alpha}}+1)^{2},\qquad\ell\neq k,\\ \mathrm{Tr}(H_{\alpha,k}H_{\beta,\ell})&=0,\qquad\beta\neq\alpha.\end{split}$ (14)
we calculate
$p_{\alpha,k}=\mathrm{Tr}(\rho P_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\mathrm{Tr}(\rho H_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\sum_{\ell=1}^{M_{\alpha}}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=\frac{a_{\alpha}}{d}+\tau(\sqrt{M_{\alpha}}+1)^{2}(M_{\alpha}r_{\alpha,k}-r_{\alpha}),$ (15)
where $r_{\alpha}=\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}$. Now, the sum of squares
$\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}+\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (16)
Remembering that $\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=\gamma_{\alpha}$, one finds
$\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}\leq\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq\gamma_{\alpha}^{2},$ (17)
which shows that the second term in eq. (16) is positive. Now, we compute the sum over an incomplete set of indices $\alpha=1,\ldots,L$;
$\begin{split}\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}&=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).\end{split}$ (18)
On the other hand, note that
$\sum_{k,\ell=1}^{M_{\alpha}}r_{\alpha,k}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right),$ (19)
and hence the purity of the state $\rho$ is equal to
$\mathrm{Tr}(\rho^{2})=\frac{1}{d}+\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (20)
After comparing eqs. (18) and (20), it becomes evident that the relation between the sum of squared $p_{\alpha,k}$ and the purity of $\rho$ is linear provided that $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, which is the condition for a conical 2-design. In this case,
$\begin{split}\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}&=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+S\sum_{\alpha=1}^{L}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &\leq\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+S\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &=S\left[\mathrm{Tr}(\rho^{2})-\frac{1}{d}\right]+\mu_{L},\end{split}$ (21)
where $\mu_{L}=\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}/d^{2}=(1/d)\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}$. For $L=N$, this result reproduces the index of coincidence $\mathcal{C}$, for which the equality holds GEAM . ∎
Observe that eq. (12) is upper bounded by its value on pure states,
$\mathcal{C}_{L}\leq\widetilde{\mathcal{C}}_{L}=\frac{d-1}{d}S+\mu_{L}.$ (22)
2 Positive maps and entanglement witnesses
Using generalized equiangular measurements that are conical 2-designs, we construct $N$ maps
$\Phi_{\alpha}[X]=\sum_{k,\ell=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}P_{\alpha,k}\mathrm{Tr}(XP_{\alpha,\ell}),$ (23)
where $\mathcal{O}^{(\alpha)}$ are orthogonal rotation matrices preserving the maximally mixed vector $\mathbf{n}_{\ast}=(1,\ldots,1)/\sqrt{d}$. Note that
$\mathrm{Tr}(\Phi_{\alpha}[X])=a_{\alpha}\gamma_{\alpha}\mathrm{Tr}(X),$ (24)
which means that $\Phi_{\alpha}$ are not trace preserving. This is important in proving the following proposition, where trace-preserving $\Phi_{\alpha}$ would impose additional conditions on $P_{\alpha,k}$.
Proposition 3.
The linear map
$\Phi=A\Phi_{0}+\sum_{\alpha=L+1}^{K}\Phi_{\alpha}-\sum_{\alpha=1}^{K}\Phi_{\alpha}$ (25)
with $A=(d-1)S-d(\mu_{K}-2\mu_{L})$ and $\Phi_{0}[X]=\mathbb{I}_{d}\mathrm{Tr}(X)/d$ is positive, where $1\leq L\leq K\leq N$.
Proof.
The condition for positivity of $\Phi$ follows directly from Mehta’s theorem Mehta
$\frac{\mathrm{Tr}(\Phi[P])^{2}}{[\mathrm{Tr}(\Phi[P])]^{2}}\leq\frac{1}{d-1}.$ (26)
It is enough to prove the above inequality for rank-1 projectors $P$ MUBs . The denominator simply reads
$[\mathrm{Tr}(\Phi[P])]^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}.$ (27)
Computing the numerator requires more attention. First, note that
$\begin{split}\mathrm{Tr}(\Phi[P])^{2}=\mathrm{Tr}\Bigg{\{}&A^{2}\Phi_{0}[P]^{2}+\sum_{\alpha,\beta=1}^{L}\Phi_{\alpha}[P]\Phi_{\beta}[P]+\sum_{\alpha,\beta=L+1}^{K}\Phi_{\alpha}[P]\Phi_{\beta}[P]+2A\sum_{\alpha=L+1}^{K}\Phi_{0}[P]\Phi_{\alpha}[P]\\ &-2A\sum_{\alpha=1}^{L}\Phi_{0}[P]\Phi_{\alpha}[P]-2\sum_{\alpha=L+1}^{K}\sum_{\beta=1}^{L}\Phi_{\alpha}[P]\Phi_{\beta}[P]\Bigg{\}}.\end{split}$ (28)
Next, using the properties of the rotation matrix
$\sum_{k=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}=\sum_{\ell=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}=1,\qquad\sum_{k=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}\mathcal{O}^{(\alpha)}_{km}=\delta_{\ell m},$ (29)
as well as the properties of $\Phi_{\alpha}$,
$\mathrm{Tr}(\Phi_{0}[P]^{2})=\frac{1}{d},\qquad\mathrm{Tr}(\Phi_{0}[P]\Phi_{\alpha}[P])=\frac{a_{\alpha}\gamma_{\alpha}}{d},\qquad\mathrm{Tr}(\Phi_{\alpha}[P]\Phi_{\beta}[P])=\frac{a_{\alpha}a_{\beta}\gamma_{\alpha}\gamma_{\beta}}{d},\quad\alpha\neq\beta,$ (30)
we simplify eq. (28) to
$\begin{split}d\mathrm{Tr}(\Phi[P])^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}-\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}+d\sum_{\alpha=1}^{K}\mathrm{Tr}(\Phi_{\alpha}[P])^{2},\end{split}$ (31)
where $\mu_{L}=\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}/d$, just as defined in Proposition 1. Now, observe that
$\mathrm{Tr}(\Phi_{\alpha}[P]^{2})=a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}$ (32)
with $p_{\alpha,k}=\mathrm{Tr}(PP_{\alpha,k})$ being calculated on a pure state $P$. In this step, we assume that $P_{\alpha,k}$ form a conical 2-design (i.e., $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$) to apply the results from eq. (22). For pure states $\rho=P$, the purity $\mathrm{Tr}(\rho^{2})=1$, and hence
$\sum_{\alpha=1}^{K}\mathrm{Tr}(\Phi_{\alpha}[P]^{2})=\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+S\sum_{\alpha=1}^{K}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+\frac{d-1}{d}S^{2}+S\mu_{K}.$ (33)
This way, eq. (34) reduces to
$d\mathrm{Tr}(\Phi[P])^{2}\leq(A+d\mu_{K}-2d\mu_{L})^{2}-\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}(1-dc_{\alpha})+S[(d-1)S+d\mu_{K}].$ (34)
Finally, note that
$S=a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=a_{\alpha}\gamma_{\alpha}(1-dc_{\alpha}),$ (35)
and therefore
$d\mathrm{Tr}(\Phi[P])^{2}\leq(A+d\mu_{K}-2d\mu_{L})^{2}+(d-1)S^{2}.$ (36)
Remembering that $A=(d-1)S-d(\mu_{K}-2\mu_{L})$, we arrive at
$d\mathrm{Tr}(\Phi[P])^{2}\leq d(d-1)S^{2}=\frac{d[\mathrm{Tr}(\Phi[P])]^{2}}{d-1},$ (37)
which proves that $\Phi$ is indeed positive. ∎
Corollary 1.
Actually, to define the positive map $\Phi$, it is enough to know $\widetilde{C}_{K}$ and $\widetilde{C}_{L}$ due to
$A=d(2\widetilde{C}_{L}-\widetilde{C}_{K}).$ (38)
The Choi-Jamiołkowski isomorphism Choi ; Jamiolkowski
$W=\sum_{k,\ell=0}^{d-1}|k\rangle\langle\ell|\otimes\Phi[|k\rangle\langle\ell|]$ (39)
establishes a one-to-one correspondence between a positive map $\Phi$ and an entanglement witness $W$, using an operational basis $|k\rangle$ in $\mathbb{C}^{d}$. The witness corresponding to $\Phi$ from Proposition 2 reads
$W=\frac{A}{d}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\sum_{\alpha=L+1}^{K}J_{\alpha}-\sum_{\alpha=1}^{K}J_{\alpha}$ (40)
where
$J_{\alpha}=\sum_{k,\ell=1}^{M_{\alpha}}\mathcal{O}_{k\ell}^{(\alpha)}\overline{P}_{\alpha,\ell}\otimes P_{\alpha,k}.$ (41)
3 Separability criteria
Take a bipartite composite system with subsystems labeled by $A$ and $B$. On each subsystem, define a conical 2-design that is a generalized equiangular measurement: $\{P_{\alpha,k}^{A}:\,k=1,\ldots,M_{\alpha}^{A};\,\alpha=1,\ldots,N_{A}\}$ and analogically for $B$. To measure correlations between subsystems, one introduces the correlation matrix with elements
$\mathcal{P}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big{[}\rho(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big{]}.$ (42)
Now, separability conditions follow in terms of its trace or trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$.
Proposition 4.
If a bipartite state $\rho$ is separable, then $\|\mathcal{P}\|_{\mathrm{Tr}}\leq\sqrt{C_{A}C_{B}}$. If additionally $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$, then the necessary separability condition reads $\mathrm{Tr}\mathcal{P}\leq\frac{C_{A}+C_{B}}{2}$. Here, $C_{A/B}$ are the indices of coincidence corresponding to conical 2-designs $\{P_{\alpha,k}^{A/B}\}$.
Proof.
For a product state $\rho=\rho_{A}\otimes\rho_{B}$, one has $\mathcal{P}_{\alpha,k;\beta,\ell}=p_{\alpha,k}^{A}p_{\beta,\ell}^{B}$ with $p_{\alpha,k}^{A}=\mathrm{Tr}(P_{\alpha,k}^{A}\rho_{A})$ and $p_{\alpha,k}^{B}=\mathrm{Tr}(P_{\alpha,k}^{B}\rho_{B})$. Therefore, using the methods from ref. ESIC , it follows that
$\|\mathcal{P}\|_{{\rm{tr}}}=\mathrm{Tr}\sqrt{\mathcal{P}^{2}}=\sqrt{\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}(p_{\alpha,k}^{A})^{2}}\sqrt{\sum_{\beta=1}^{N_{B}}\sum_{\ell=1}^{M_{\beta}^{B}}(p_{\beta,\ell}^{B})^{2}}=\sqrt{C_{A}C_{B}}.$ (43)
Note that the trace of $\mathcal{P}$ is well defined only under the additional condition that $\mathcal{P}$ is a square matrix. This is the case as long as $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$. Then,
$\mathrm{Tr}\mathcal{P}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{A}p_{\alpha,k}^{B}\leq\frac{1}{2}\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}\Big{[}(p_{\alpha,k}^{A})^{2}+(p_{\alpha,k}^{B})^{2}\Big{]}=\frac{C_{A}+C_{B}}{2}.$ (44)
Finally, from the convexity of the trace norm and linearity of the trace, the above conditions in terms of $\mathrm{Tr}\mathcal{P}$ and $\|\mathcal{P}\|_{\mathrm{Tr}}$ extend from product states to all separable states. ∎
The above results include as special cases the separability criteria for mutually unbiased bases Spengler and measurements ChenMa , general SIC POVMs ChenLi , $(N,M)$-POVMs SIC-MUB , generalized symmetric measurements SIC-MUB_general , and measurements based on equiangular tight frames W_ETS .
4 Acknowledgements
This research was funded in whole or in part by the National Science Centre, Poland, Grant number 2021/43/D/ST2/00102. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.

### v3

@@
 Additionally, we introduce separability criteria based on the correlation matrix and analyze them for various types of measurements.
+1 Introduction
+Entanglement is one of the most intriguing and fundamental features of quantum mechanics, playing a central role in quantum information science, including quantum computation, quantum cryptography, and teleportation protocols HHHH ; Nielsen .
+However, detecting whether a given quantum state is entangled or separable can be a highly nontrivial task, especially in higher-dimensional or multipartite systems Gurvits .
+The Peres-Horodecki criterion, also known as the PPT (positive partial transpose) criterion, is a widely used method for detecting quantum entanglement in bipartite systems.
+It was introduced by Peres PhysRevLett.77.1413 and later fully characterized by the Horodecki family PHHH ; HORODECKI1997333 .
+This criterion provides a necessary and sufficient separability condition for qubit-qubit and qubit-qutrit systems.
+In other instances, only some PPT states are separable.
+The rest are PPT entangled states, describing bound entanglement HHH_bound .
+Therefore, more refined detection methods are needed.
+The CCNR (computable cross-norm or realignment) criterion Rudolph ; HHH_OSID ; ChenWu ; ChenWu2 rearranges the density matrix via realignment and computes its trace norm.
+The nonlinear entanglement criteria use nonlinear functionals of the density matrix, like the purity or Tsallis entropy GuhneLev .
+The correlation matrix criterion deVicente ; deVicente2 and the covariance matrix criterion CMM1 ; Gittsovich rely on calculations upon matrices constructed for given states.
+Entanglement witnesses provide a powerful and practical tool for identifying entangled states.
+A Hermitian operator is called an entanglement witness if it is positive on all separable states and negative on at least one entangled state.
+While a single entanglement witness cannot detect all entangled states, every entangled state can, in principle, be detected by some witness Terhal1 ; Terhal2 .
+Importantly, entanglement detection can be carried out without requiring full knowledge of the state.
+Thus, entanglement witnesses form a central component of experimental and theoretical strategies for detecting quantum entanglement Barbieri ; Lewenstein .
+In the construction of entanglement witnesses, one often uses positive but not completely positive maps defined via measurement operators.
+A wide class of witnesses from mutually unbiased bases (MUBs) Durt was derived in ref.
+MUBs and analyzed in more details for qutrit-qutrit systems.
+Further generalizations of this formalism include replacing the mutually unbiased bases with mutually unbiased measurements (MUMs) Li ; P_maps , symmetric informationally complete measurements (SIC POVMs) EW-SIC , symmetric measurements SIC-MUB , and equiangular tight frames W_ETS .
+Witnesses for composite systems of unequal dimensions were also considered EW-2MUB .
+The number of measurement operators required to detect bound entanglement was analyzed in refs.
+bound_ent ; MUM_purity .
+In this paper, we provide a construction method for positive (but not trace preserving) maps and the associated entanglement witnesses from informationally overcomplete measurements.
+For this purpose, we apply the generalized equiangular measurements (GEAMs) GEAM , which are sets of mutually unbiased non-projective equiangular tight frames.
+Importantly, we focus on those measurements that form conical 2-designs, because only then a partial index of coincidence can be derived.
+Interestingly, the entanglement witnesses from GEAMs are parametrized only by two partial indices of coincidence.
+Finally, as an alternative way to detect entanglement, we provide enhanced separability criteria via the correlation matrix.
 2 Measurements from generalized equiangular tight frames
@@
 Definition 1.
-(GEAM ) A collection of $N$ generalized equiangular lines $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}$ is a generalized equiangular measurement provided that
+(GEAM ) A collection of $N$ generalized equiangular lines $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}$ is a generalized equiangular measurement (GEAM) provided that $\sum_{\alpha=1}^{N}M_{\alpha}=d^{2}+N-1$ and
 $\begin{split}\mathrm{Tr}(P_{\alpha,k})&=a_{\alpha},\\ \mathrm{Tr}(P_{\alpha,k}^{2})&=b_{\alpha}\mathrm{Tr}(P_{\alpha,k})^{2},\\ \mathrm{Tr}(P_{\alpha,k}P_{\alpha,\ell})&=c_{\alpha}\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\alpha,\ell}),\qquad k\neq\ell,\\ \mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})&=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell}),\qquad\alpha\neq\beta,\end{split}$ (1)
@@
 $\frac{1}{d}<b_{\alpha}\leq\frac{1}{d}\min\{d,M_{\alpha}\}.$ (3)
-Generalized equiangular measurements can be constructed from Hermitian orthonormal bases $\{G_{0}=\mathbb{I}_{d}/\sqrt{d},G_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ with traceless $G_{\alpha,k}$.
+Generalized equiangular measurements can be constructed from Hermitian orthonormal bases $\{G_{0}=\mathbb{I}_{d}/\sqrt{d},G_{\alpha,k}:\,k=1,\ldots,M_{\alpha}-1;\,\alpha=1,\ldots,N\}$ with traceless operators $G_{\alpha,k}$.
 Namely, one has GEAM
 $P_{\alpha,k}=\frac{a_{\alpha}}{d}\mathbb{I}_{d}+\tau_{\alpha}H_{\alpha,k},$ (4)
-where the parameter
-$\tau_{\alpha}^{2}=\frac{a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})}{M_{\alpha}(\sqrt{M_{\alpha}}+1)^{2}}$ (5)
-is chosen this way to guarantee $P_{\alpha,k}\geq 0$. The traceless operators $H_{\alpha,k}$ are related to $G_{\alpha,k}$ in the following way,
+where the traceless operators $H_{\alpha,k}$ are related to $G_{\alpha,k}$ in the following way,
 $H_{\alpha,k}=\left\{\begin{aligned} &G_{\alpha}-\sqrt{M_{\alpha}}(1+\sqrt{M_{\alpha}})G_{\alpha,k},\quad k=1,\ldots,M_{\alpha}-1,\\ &(1+\sqrt{M_{\alpha}})G_{\alpha},\qquad k=M_{\alpha},\end{aligned}\right.$ (5)
-where $G_{\alpha}=\sum_{k=1}^{M_{\alpha}-1}G_{\alpha,k}$.
+with $G_{\alpha}=\sum_{k=1}^{M_{\alpha}-1}G_{\alpha,k}$. The real parameter
+$\tau_{\alpha}=\pm\sqrt{\frac{a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})}{M_{\alpha}(\sqrt{M_{\alpha}}+1)^{2}}}$ (6)
+is chosen to guarantee that $P_{\alpha,k}\geq 0$.
 There exists an important class of generalized equiangular measurements that are conical 2-designs Graydon ; Graydon2 – that is, they satisfy
 $\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}\otimes P_{\alpha,k}=\kappa_{+}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\kappa_{-}\mathbb{F}_{d}$ (7)
-for real parameters $\kappa_{+}\geq\kappa_{-}>0$, where $\mathbb{F}_{d}=\sum_{m,n=1}^{d}|m\rangle\langle n|\otimes|n\rangle\langle m|$ is the flip operator.
-It has been shown in ref.
-GEAM that a linear dependence between purity $\mathrm{Tr}(\rho^{2})$ of a state $\rho$ and the corresponding index of coincidence Rastegin5
+for real parameters $\kappa_{+}\geq\kappa_{-}>0$, where $\mathbb{F}_{d}=\sum_{m,n=0}^{d-1}|m\rangle\langle n|\otimes|n\rangle\langle m|$ is the flip operator.
+This equation holds as long as the GEAM parameters satisfy $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$ with
+$0<S\leq\min\left\{\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\right\}.$ (8)
+In this case, $\kappa_{\pm}$ in eq. (7) are given by GEAM
+$\kappa_{+}=\mu_{N}-\frac{S}{d},\qquad\kappa_{-}=S,\qquad\mu_{N}=\frac{1}{d}\sum_{\alpha=1}^{N}a_{\alpha}\gamma_{\alpha}.$ (9)
+It has been shown that a linear dependence between the purity $\mathrm{Tr}(\rho^{2})$ of a state $\rho$ and its index of coincidence Rastegin5
 $\mathcal{C}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad p_{\alpha,k}=\mathrm{Tr}(P_{\alpha,k}\rho),$ (10)
@@
 Proposition 1.
-The generalized equiangular measurement whose parameters satisfy $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, where
-$0<S\leq\min\left\{\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\right\},$ (9)
-is a conical 2-design with real parameters
-$\kappa_{+}=\mu-\frac{S}{d},\qquad\kappa_{-}=S,\qquad\mu=\frac{1}{d}\sum_{\alpha=1}^{N}a_{\alpha}\gamma_{\alpha}.$ (10)
-Observe that each generalized equiangular tight frame has its own index of coincidence (after rescaling)
+(GEAM ] If the generalized equiangular measurement is a conical 2-design, then the associated index of coincidence satisfies
+$\mathcal{C}=S\left(\mathrm{Tr}\rho^{2}-\frac{1}{d}\right)+\mu\leq\frac{d-1}{d}S+\mu,$ (11)
+where $\mu=(1/d)\sum_{\alpha=1}^{N}a_{\alpha}\gamma_{\alpha}$.
+Interestingly, the above proposition can be further generalized. Observe that each generalized equiangular tight frame has its own index of coincidence (after rescaling)
 $\mathcal{C}_{\alpha}=\frac{1}{\gamma_{\alpha}^{2}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad\frac{1}{\gamma_{\alpha}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=1.$ (12)
-Therefore, it is possible to derive an inequality for sums of ssquared probabilities, similar to that for few mutually unbiased bases Wu .
+Therefore, it is possible to derive an inequality for partial sums of squared probabilities, similar to that for few mutually unbiased bases Wu .
 Proposition 2.
-If $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a conical 2-design, then
+If $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a conical 2-design, then the partial index of coincidence
 $\mathcal{C}_{L}=\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq S\left(\mathrm{Tr}\rho^{2}-\frac{1}{d}\right)+\mu_{L},$ (13)
@@
 $\begin{split}\mathrm{Tr}(H_{\alpha,k}^{2})&=(M_{\alpha}-1)(\sqrt{M_{\alpha}}+1)^{2},\\ \mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})&=-(\sqrt{M_{\alpha}}+1)^{2},\qquad\ell\neq k,\\ \mathrm{Tr}(H_{\alpha,k}H_{\beta,\ell})&=0,\qquad\beta\neq\alpha.\end{split}$ (15)
-we calculate
+we calculate the probabilities
 $p_{\alpha,k}=\mathrm{Tr}(\rho P_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\mathrm{Tr}(\rho H_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\sum_{\ell=1}^{M_{\alpha}}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=\frac{a_{\alpha}}{d}+\tau(\sqrt{M_{\alpha}}+1)^{2}(M_{\alpha}r_{\alpha,k}-r_{\alpha}),$ (16)
-where $r_{\alpha}=\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}$. Now, the sum of squares
+where $r_{\alpha}=\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}$. Now, the sum of squares over the second index satisfies
 $\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}+\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (17)
 Remembering that $\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=\gamma_{\alpha}$, one finds
-$\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}\leq\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq\gamma_{\alpha}^{2},$ (17)
-which shows that the second term in eq. (16) is positive. Now, we compute the sum over an incomplete set of indices $\alpha=1,\ldots,L$;
+$\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\geq\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}},$ (18)
+where we used the fact that $\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}$ reaches its lower bound when all $p_{\alpha,k}=\gamma_{\alpha}/M_{\alpha}$.
+This shows that the second term on the right hand-side of eq.
+(17) is positive.
+Now, we compute the sum over an incomplete set of indices $\alpha=1,\ldots,L$;
 $\begin{split}\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}&=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).\end{split}$ (19)
-On the other hand, note that
+On the other hand, from eq. (14), we find the purity of the state $\rho$ is equal to
+$\mathrm{Tr}(\rho^{2})=\frac{1}{d}+\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right),$ (20)
+where we used properties (15) to compute
 $\sum_{k,\ell=1}^{M_{\alpha}}r_{\alpha,k}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right),$ (21)
-and hence the purity of the state $\rho$ is equal to
-$\mathrm{Tr}(\rho^{2})=\frac{1}{d}+\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (20)
 After comparing eqs.
-(18) and (20), it becomes evident that the relation between the sum of squared $p_{\alpha,k}$ and the purity of $\rho$ is linear provided that $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, which is the condition for a conical 2-design.
+(19) and (20), it becomes evident that the relation between the sum of squared $p_{\alpha,k}$ and the purity of $\rho$ is linear provided that $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, which is exactly the condition for a conical 2-design from Proposition 1.
 In this case,
@@
 Observe that eq. (13) is upper bounded by its value on pure states,
-$\mathcal{C}_{L}\leq\widetilde{\mathcal{C}}_{L}=\frac{d-1}{d}S+\mu_{L}.$ (22)
+$\mathcal{C}_{L}\leq\widetilde{\mathcal{C}}_{L}=\frac{d-1}{d}S+\mu_{L},$ (23)
+where $\mathrm{Tr}(\rho^{2})=1$.
 3 Positive maps and entanglement witnesses
@@
 which means that $\Phi_{\alpha}$ are not trace preserving.
-This is important in proving the following proposition, where trace-preserving $\Phi_{\alpha}$ would impose additional conditions on $P_{\alpha,k}$.
+This is important in proving the following proposition, where trace-preserving $\Phi_{\alpha}$ would result in imposing additional conditions on $P_{\alpha,k}$.
 Proposition 3.
 The linear map
-$\Phi=A\Phi_{0}+\sum_{\alpha=L+1}^{K}\Phi_{\alpha}-\sum_{\alpha=1}^{K}\Phi_{\alpha}$ (25)
-with $A=(d-1)S-d(\mu_{K}-2\mu_{L})$ and $\Phi_{0}[X]=\mathbb{I}_{d}\mathrm{Tr}(X)/d$ is positive, where $1\leq L\leq K\leq N$.
+$\Phi=A\Phi_{0}+\sum_{\alpha=L+1}^{K}\Phi_{\alpha}-\sum_{\alpha=1}^{L}\Phi_{\alpha}$ (26)
+with $A=d(2\widetilde{\mathcal{C}}_{L}-\widetilde{\mathcal{C}}_{K})$ and $\Phi_{0}[X]=\mathbb{I}_{d}\mathrm{Tr}(X)/d$ is positive, where $1\leq L\leq K\leq N$.
 Proof.
@@
 with $p_{\alpha,k}=\mathrm{Tr}(PP_{\alpha,k})$ being calculated on a pure state $P$.
-In this step, we assume that $P_{\alpha,k}$ form a conical 2-design (i.e., $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$) to apply the results from eq.
+In this step, it becomes necessary to assume that $P_{\alpha,k}$ form a conical 2-design (i.e., $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$).
+Then, we can apply the results from eq.
 (23).
@@
 $d\mathrm{Tr}(\Phi[P])^{2}\leq(A+d\mu_{K}-2d\mu_{L})^{2}+(d-1)S^{2}.$ (37)
-Remembering that $A=(d-1)S-d(\mu_{K}-2\mu_{L})$, we arrive at
+Now, if we recall the definition of $A$ and $\widetilde{\mathcal{C}}_{L}$, we see that
+$A=d(2\widetilde{\mathcal{C}}_{L}-\widetilde{\mathcal{C}}_{K})=(d-1)S-d(\mu_{K}-2\mu_{L}),$ (38)
+and eq. (37) further simplifies to
 $d\mathrm{Tr}(\Phi[P])^{2}\leq d(d-1)S^{2}=\frac{d[\mathrm{Tr}(\Phi[P])]^{2}}{d-1},$ (39)
-which proves that $\Phi$ is indeed positive. ∎
-Corollary 1.
-Actually, to define the positive map $\Phi$, it is enough to know $\widetilde{C}_{K}$ and $\widetilde{C}_{L}$ due to
-$A=d(2\widetilde{C}_{L}-\widetilde{C}_{K}).$ (38)
-The Choi-Jamiołkowski isomorphism Choi ; Jamiolkowski
-$W=\sum_{k,\ell=0}^{d-1}|k\rangle\langle\ell|\otimes\Phi[|k\rangle\langle\ell|]$ (39)
-establishes a one-to-one correspondence between a positive map $\Phi$ and an entanglement witness $W$, using an operational basis $|k\rangle$ in $\mathbb{C}^{d}$.
-The witness corresponding to $\Phi$ from Proposition 2 reads
-$W=\frac{A}{d}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\sum_{\alpha=L+1}^{K}J_{\alpha}-\sum_{\alpha=1}^{K}J_{\alpha}$ (40)
+which proves that $\Phi$ is indeed a positive map. ∎
+Positive but not completely positive maps are used in the theory of quantum entanglement to construct block-positive operators that detect entangled (non-separable) states.
+Through the Choi-Jamiołkowski isomorphism Choi ; Jamiolkowski
+$W=\sum_{m,n=0}^{d-1}|m\rangle\langle n|\otimes\Phi[|m\rangle\langle n|],$ (40)
+a one-to-one correspondence is established between a positive map $\Phi$ and an entanglement witness $W$, using an operational basis $|k\rangle$ in $\mathbb{C}^{d}$.
+In particular, the witness corresponding to $\Phi$ from Proposition 2 reads
+$W=\Big{(}2\widetilde{\mathcal{C}}_{L}-\widetilde{\mathcal{C}}_{K}\Big{)}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\sum_{\alpha=L+1}^{K}J_{\alpha}-\sum_{\alpha=1}^{L}J_{\alpha}$ (41)
 where
 $J_{\alpha}=\sum_{k,\ell=1}^{M_{\alpha}}\mathcal{O}_{k\ell}^{(\alpha)}\overline{P}_{\alpha,\ell}\otimes P_{\alpha,k}.$ (42)
+Note that $W$ is parametrized only by two partial indices of coincidence: one associated with the number of negative terms $L$ and the other with the number of all terms $K$.
+Example 1.
+A trivial example follows from the choice $\mathcal{O}^{(\alpha)}=\mathbb{I}_{M_{\alpha}}$ and $L=K=N$. Then, using eq. (7), one gets
+$W=S\left(\mathbb{I}_{d^{2}}-dP_{+}\right),$ (43)
+where $P_{+}=(1/d)\sum_{m,n=0}^{d-1}|m\rangle\langle n|\otimes|m\rangle\langle n|$ is a maximally entangled state. This is the entanglement witness for the reduction map EW_reduction ; P_maps .
 4 Separability criteria
-Take a bipartite composite system with subsystems labeled by $A$ and $B$.
-On each subsystem, define a conical 2-design that is a generalized equiangular measurement: $\{P_{\alpha,k}^{A}:\,k=1,\ldots,M_{\alpha}^{A};\,\alpha=1,\ldots,N_{A}\}$ and analogically for $B$.
-To measure correlations between subsystems, one introduces the correlation matrix with elements
+Take a bipartite composite system with subsystems labeled by $A$ and $B$, so that the total state $\rho:\mathcal{H}_{A}\otimes\mathcal{H}_{B}\to\mathcal{H}_{A}\otimes\mathcal{H}_{B}$.
+On each subsystem, define a generalized equiangular measurement: $\{P_{\alpha,k}^{A}:\,k=1,\ldots,M_{\alpha}^{A};\,\alpha=1,\ldots,N_{A}\}$ and analogically for $B$.
+No further assumptions are made regarding the measurement parameters $(a_{\alpha},b_{\alpha},c_{\alpha},f)$, so in general they can depend on the choice of subsystem.
+Now, to measure correlations between subsystems $A$ and $B$, one introduces the correlation matrix deVicente ; deVicente2 with elements
 $\mathcal{P}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big{[}\rho(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big{]}.$ (44)
-Now, separability conditions follow in terms of its trace or trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$.
+The separability conditions follow in terms of its trace or trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$, and as it turns out they depend on the maximal indices of coincidence $\widetilde{\mathcal{C}}^{A}$, $\widetilde{\mathcal{C}}^{B}$ associated with $\{P_{\alpha,k}^{A}\}$ and $\{P_{\alpha,k}^{B}\}$, respectively.
 Proposition 4.
-If a bipartite state $\rho$ is separable, then $\|\mathcal{P}\|_{\mathrm{Tr}}\leq\sqrt{C_{A}C_{B}}$.
-If additionally $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$, then the necessary separability condition reads $\mathrm{Tr}\mathcal{P}\leq\frac{C_{A}+C_{B}}{2}$.
-Here, $C_{A/B}$ are the indices of coincidence corresponding to conical 2-designs $\{P_{\alpha,k}^{A/B}\}$.
+If a bipartite state $\rho$ is separable, then the necessary separability condition reads
+$\mathrm{Tr}\mathcal{P}\leq\frac{\widetilde{\mathcal{C}}^{A}+\widetilde{\mathcal{C}}^{B}}{2},$ (45)
+where the GEAMs are chosen in such a way that $N_{A}=N_{B}$ and $M_{\alpha}^{A}=M_{\alpha}^{B}$.
 Proof.
 For a product state $\rho=\rho_{A}\otimes\rho_{B}$, one has $\mathcal{P}_{\alpha,k;\beta,\ell}=p_{\alpha,k}^{A}p_{\beta,\ell}^{B}$ with $p_{\alpha,k}^{A}=\mathrm{Tr}(P_{\alpha,k}^{A}\rho_{A})$ and $p_{\alpha,k}^{B}=\mathrm{Tr}(P_{\alpha,k}^{B}\rho_{B})$.
-Therefore, using the methods from ref.
-ESIC , it follows that
-$\|\mathcal{P}\|_{{\rm{tr}}}=\mathrm{Tr}\sqrt{\mathcal{P}^{2}}=\sqrt{\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}(p_{\alpha,k}^{A})^{2}}\sqrt{\sum_{\beta=1}^{N_{B}}\sum_{\ell=1}^{M_{\beta}^{B}}(p_{\beta,\ell}^{B})^{2}}=\sqrt{C_{A}C_{B}}.$ (43)
 Note that the trace of $\mathcal{P}$ is well defined only under the additional condition that $\mathcal{P}$ is a square matrix.
-This is the case as long as $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$.
-Then,
-$\mathrm{Tr}\mathcal{P}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{A}p_{\alpha,k}^{B}\leq\frac{1}{2}\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}\Big{[}(p_{\alpha,k}^{A})^{2}+(p_{\alpha,k}^{B})^{2}\Big{]}=\frac{C_{A}+C_{B}}{2}.$ (44)
-Finally, from the convexity of the trace norm and linearity of the trace, the above conditions in terms of $\mathrm{Tr}\mathcal{P}$ and $\|\mathcal{P}\|_{\mathrm{Tr}}$ extend from product states to all separable states.
+This is the case as long as the GEAMs on both subsystems are of equal number: $N_{A}=N_{B}\equiv N$ and $M_{\alpha}^{A}=M_{\alpha}^{B}\equiv M_{\alpha}$.
+Therefore, under this additional condition,
+$\mathrm{Tr}\mathcal{P}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{A}p_{\alpha,k}^{B}\leq\frac{1}{2}\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}\Big{[}(p_{\alpha,k}^{A})^{2}+(p_{\alpha,k}^{B})^{2}\Big{]}=\frac{\mathcal{C}^{A}+\mathcal{C}^{B}}{2}\leq\frac{\widetilde{\mathcal{C}}^{A}+\widetilde{\mathcal{C}}^{B}}{2}.$ (46)
+From the linearity of the trace, this result extends from product states to all separable states. ∎
+Proposition 5.
+If a bipartite state $\rho$ is separable, then $\|\mathcal{P}\|_{\mathrm{Tr}}\leq\sqrt{\widetilde{\mathcal{C}}^{A}\widetilde{\mathcal{C}}^{B}}$.
+Proof.
+Using the notation from the proof to Proposition 4 and the methods from ref. ESIC , we show that
+$\|\mathcal{P}\|_{{\rm{tr}}}=\mathrm{Tr}\sqrt{\mathcal{P}^{2}}=\sqrt{\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}(p_{\alpha,k}^{A})^{2}}\sqrt{\sum_{\beta=1}^{N_{B}}\sum_{\ell=1}^{M_{\beta}^{B}}(p_{\beta,\ell}^{B})^{2}}=\sqrt{\mathcal{C}^{A}\mathcal{C}^{B}}\leq\sqrt{\widetilde{\mathcal{C}}^{A}\widetilde{\mathcal{C}}^{B}}.$ (47)
+Finally, from the convexity of the trace norm, this condition extends from product states to all separable states. ∎
+The above results include as special cases the separability criteria for mutually unbiased bases Spengler ; Liu and measurements ChenMa ; Liu2 , general SIC POVMs ChenLi , $(N,M)$-POVMs SIC-MUB , generalized symmetric measurements SIC-MUB_general , and measurements based on equiangular tight frames W_ETS .
+An improved but more computationally demanding criterion is based on the matrix
+$\mathcal{R}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big{[}(\rho-\rho_{A}\otimes\rho_{B})(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big{]},$ (48)
+where $\rho^{A}=\mathrm{Tr}_{B}(\rho)$ and $\rho^{B}=\mathrm{Tr}_{A}(\rho)$ are the marginal states of $\rho$, obtained after calculating the partial trace over one subsystem.
+Proposition 6.
+If a bipartite state $\rho$ is separable, then
+$\|\mathcal{R}\|_{{\rm{tr}}}=\sqrt{\widetilde{\mathcal{C}}^{A}-\mathcal{C}^{A}(\rho^{A})}\sqrt{\widetilde{\mathcal{C}}^{B}-\mathcal{C}^{B}(\rho^{B})},$ (49)
+where $\mathcal{C}^{A}(\rho^{A})$ and $\mathcal{C}^{B}(\rho^{B})$ are indices of coincidence for the marginal states of $\rho$.
+Proof.
+Any separable state $\rho$ can be written as
+$\rho=\sum_{j=1}^{n}q_{j}\rho^{A}_{j}\otimes\rho^{B}_{j},$ (50)
+where $q_{j}$ is a probability distribution. The corresponding marginal states read
+$\rho^{A}=\sum_{j=1}^{n}q_{j}\rho^{A}_{j},\qquad\rho^{B}=\sum_{j=1}^{n}q_{j}\rho^{B}_{j}.$ (51)
+Therefore, one finds
+$(\rho-\rho^{A}\otimes\rho^{B})(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})=\sum_{j=1}^{n}q_{j}\left(\rho^{A}_{j}-\rho^{A}\right)P_{\alpha,k}^{A}\otimes\left(\rho^{B}_{j}-\rho^{B}\right)P_{\beta,\ell}^{B}.$ (52)
+Now, introduce two vectors, $v_{j}^{A}$ and $v_{j}^{B}$, with the following elements,
+$\left(v^{A}_{j}\right)_{\alpha,k}=\mathrm{Tr}\left[\left(\rho^{A}_{j}-\rho^{A}\right)P_{\alpha,k}^{A}\right],\qquad\left(v^{B}_{j}\right)_{\beta,\ell}=\mathrm{Tr}\left[\left(\rho^{B}_{j}-\rho^{B}\right)P_{\beta,\ell}^{B}\right].$ (53)
+This allows us to simplify the notation after taking the trace of eq. (52),
+$\mathcal{R}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big{[}(\rho-\rho_{A}\otimes\rho_{B})(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big{]}=\sum_{j=1}^{n}q_{j}\left(v^{A}_{j}\right)_{\alpha,k}\left(v^{B}_{j}\right)_{\beta,\ell}.$ (54)
+Next, from the triangle inequality of the trace norm, we calculate
+$\|\mathcal{R}_{\alpha,k;\beta,\ell}\|_{\rm{tr}}\leq\sum_{j=1}^{n}q_{j}\|v^{A}_{j}\|_{\rm{tr}}\|v^{B}_{j}\|_{\rm{tr}}=\sum_{j=1}^{n}q_{j}\|v^{A}_{j}\|\|v^{B}_{j}\|,$ (55)
+where we used the fact that the trace norm of a vector is simply a vector norm. From the Cauchy-Schwarz inequality,
+$\sum_{j=1}^{n}q_{j}\|v^{A}_{j}\|\|v^{B}_{j}\|\leq\sqrt{\sum_{j=1}^{n}q_{j}\|v^{A}_{j}\|^{2}}\sqrt{\sum_{m=1}^{n}q_{m}\|v^{B}_{m}\|^{2}}.$ (56)
+In analogy to eq. (10), let us denote the trace expressions by
+$p_{\alpha,k}^{A}(\rho^{A})=\mathrm{Tr}(P_{\alpha,k}^{A}\rho^{A}),\qquad p_{\beta,\ell}^{B}(\rho^{B})=\mathrm{Tr}(P_{\beta,\ell}^{B}\rho^{B}),$ (57)
+where the corresponding indices of coincidence are given by
+$\mathcal{C}^{A}(\rho^{A})=\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}\left[p_{\alpha,k}^{A}(\rho^{A})\right]^{2},\qquad\mathcal{C}^{B}(\rho^{B})=\sum_{\beta=1}^{N_{B}}\sum_{\ell=1}^{M_{\beta}^{B}}\left[p_{\beta,\ell}^{B}(\rho^{B})\right]^{2}.$ (58)
+Then, the term under the square root reduces to
+$\sum_{j=1}^{n}q_{j}\|v^{A}_{j}\|^{2}=\sum_{j=1}^{n}q_{j}\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}\Big{(}\mathrm{Tr}\left[\left(\rho^{A}_{j}-\rho^{A}\right)P_{\alpha,k}^{A}\right]\Big{)}^{2}=\sum_{j=1}^{n}q_{j}\mathcal{C}^{A}(\rho_{j}^{A})-\mathcal{C}^{A}(\rho^{A})\leq\widetilde{\mathcal{C}}^{A}-\mathcal{C}^{A}(\rho^{A}),$ (59)
+with $\widetilde{\mathcal{C}}^{A}$ being the upper bound of $\mathcal{C}^{A}$, and analogically for the subsystem $B$.
+After substituting these formulas into eq.
+(56) and then eq.
+(55), we reproduce eq.
+(49).
 ∎
-The above results include as special cases the separability criteria for mutually unbiased bases Spengler and measurements ChenMa , general SIC POVMs ChenLi , $(N,M)$-POVMs SIC-MUB , generalized symmetric measurements SIC-MUB_general , and measurements based on equiangular tight frames W_ETS .
+This generalizes the separability criteria for popular measurements: mutually unbiased bases, mutually unbiased measurements, general SIC POVMs Liu2 ; Shen ; Liu3 , and also for equiangular tight frames TangWu ; QiZhang .
+Moreover, our results allow to simplify these criteria significantly by proving that they depend only on the index of coincidence for the marginal state and its upper bound.
+5 Conclusions
+In this paper, we applied the recently introduced generalized equiangular measurements to quantum entanglement detection.
+It was shown that, as long as GEAMs are conical 2-designs, the purity of the state is directly related to the partial index of coincidence, where the summation of squared probabilities is performed over few generalized equiangular tight frames.
+Such measurements were then used to define a class of positive but not trace-preserving maps and the corresponding entanglement witnesses.
+Interestingly, these witnesses are parametrized solely by two partial indices of coincidence.
+Next, we used the generalized equiangular measurements to construct separability criteria for bipartite states with subsystems of arbitrary dimensions.
+In further research, it would be interesting to examine the properties of entanglement witnesses, such as extremality or optimality.
+Construction of witnesses and separability criteria from measurements that are not conical 2-designs remains as an open question.
 6 Acknowledgements

### v4

@@
 Abstract
-We use operators from generalized equiangular measurements to construct positive maps.
-Their positivity follows from the inequality for indices of coincidence corresponding to few equiangular tight frames.
+In quantum information theory, measurements are represented by positive, operator-valued measures.
+In this paper, we use the operators corresponding to generalized equiangular measurements to construct positive maps.
+Their positivity follows from the properties of index of coincidence for few equiangular tight frames.
 These maps give rise to entanglement witnesses, which include as special cases many important classes considered in the literature.
@@
 Thus, entanglement witnesses form a central component of experimental and theoretical strategies for detecting quantum entanglement Barbieri ; Lewenstein .
+Of particular interest are indecomposable entanglement witnesses, which allow one to detect PPT entangled states.
+Such states are also called bound entangled HHH_bound ; HORODECKI1997333 ; bound_review .
+They are non-distillable, which means that no protocols exist that transform them into the states that are maximally entangled Clarisse ; Horodeccy2 .
+Bound entangled states find important applications e.g.
+in quantum cryptography Ozols ; Augusiak and quantum metrology Toth ; Vertesi .
+Moreover, bound entanglement can be activated to provide resources important for quantum information and communication technologies Smolin ; Kaneda ; Ozaydin .
 In the construction of entanglement witnesses, one often uses positive but not completely positive maps defined via measurement operators.
@@
 Interestingly, the entanglement witnesses from GEAMs are parametrized only by two partial indices of coincidence.
+Examples of indecomposable witnesses (detecting bound entanglement) are presented for the qutrit-qutrit scenario.
 Finally, as an alternative way to detect entanglement, we provide enhanced separability criteria via the correlation matrix.
@@
 Definition 1.
-(GEAM ) A collection of $N$ generalized equiangular lines $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}$ is a generalized equiangular measurement (GEAM) provided that $\sum_{\alpha=1}^{N}M_{\alpha}=d^{2}+N-1$ and
+(GEAM ) A collection of $N$ generalized equiangular tight frames $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}>0$ is a generalized equiangular measurement (GEAM) provided that $\sum_{\alpha=1}^{N}M_{\alpha}=d^{2}+N-1$ and
 $\begin{split}\mathrm{Tr}(P_{\alpha,k})&=a_{\alpha},\\ \mathrm{Tr}(P_{\alpha,k}^{2})&=b_{\alpha}\mathrm{Tr}(P_{\alpha,k})^{2},\\ \mathrm{Tr}(P_{\alpha,k}P_{\alpha,\ell})&=c_{\alpha}\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\alpha,\ell}),\qquad k\neq\ell,\\ \mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})&=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell}),\qquad\alpha\neq\beta,\end{split}$ (1)
@@
 where the traceless operators $H_{\alpha,k}$ are related to $G_{\alpha,k}$ in the following way,
-$H_{\alpha,k}=\left\{\begin{aligned} &G_{\alpha}-\sqrt{M_{\alpha}}(1+\sqrt{M_{\alpha}})G_{\alpha,k},\quad k=1,\ldots,M_{\alpha}-1,\\ &(1+\sqrt{M_{\alpha}})G_{\alpha},\qquad k=M_{\alpha},\end{aligned}\right.$ (5)
+$H_{\alpha,k}=\left\{\begin{aligned} &G_{\alpha}-\sqrt{M_{\alpha}}(1+\sqrt{M_{\alpha}})G_{\alpha,k},\quad k=1,\ldots,M_{\alpha}-1,\\ &(1+\sqrt{M_{\alpha}})G_{\alpha},\qquad k=M_{\alpha},\end{aligned}\right$.
+(5)
 with $G_{\alpha}=\sum_{k=1}^{M_{\alpha}-1}G_{\alpha,k}$. The real parameter
@@
 It is enough to prove the above inequality for rank-1 projectors $P$ MUBs . The denominator simply reads
-$[\mathrm{Tr}(\Phi[P])]^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}.$ (28)
+$\mathrm{Tr}(\Phi[P])]^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}.$ (28)
 Computing the numerator requires more attention. First, note that
@@
 Note that $W$ is parametrized only by two partial indices of coincidence: one associated with the number of negative terms $L$ and the other with the number of all terms $K$.
-Example 1.
-A trivial example follows from the choice $\mathcal{O}^{(\alpha)}=\mathbb{I}_{M_{\alpha}}$ and $L=K=N$. Then, using eq. (7), one gets
+By choosing specific generalized equiangular measurement operators $P_{\alpha,k}$ and the rotation matrices $\mathcal{O}_{k\ell}^{(\alpha)}$, one finds explicit forms of entanglement witnesses.
+Note that the construction method of $\Phi$ from Proposition 3 allows also for completely positive maps.
+However, $W$ corresponding to a completely positive $\Phi$ does not detect quantum entanglement due to $\mathrm{Tr}(W\rho)\geq 0$ for all $\rho$.
+Obtaining analytical formulas for complete positivity of a general $\Phi$ is a highly non-trivial task, therefore it is crucial to check for it after constructing explicit examples.
+Proposition 4.
+A trivial case follows from the choice $\mathcal{O}^{(\alpha)}=\mathbb{I}_{M_{\alpha}}$ and $L=K=N$. Then, using eq. (7), one gets
 $W=S\left(\mathbb{I}_{d^{2}}-dP_{+}\right),$ (43)
 where $P_{+}=(1/d)\sum_{m,n=0}^{d-1}|m\rangle\langle n|\otimes|m\rangle\langle n|$ is a maximally entangled state. This is the entanglement witness for the reduction map EW_reduction ; P_maps .
+4 Detection of bound entanglement
+[…]
+In what follows, we provide qutrit ($d=3$) examples of indecomposable entanglement witnesses from generalized equiangular measurements that go beyond the well-known classes of equinumerous measurements.
+The GEAMs are constructed as in ref.
+GEAM , using the normalized Gell-Mann matrices: $G_{0}=\mathbb{I}_{3}/\sqrt{3}$ and
+$G_{1,1}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\qquad G_{1,2}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\qquad G_{2,1}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix}$$\qquad$ $G_{2,2}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}$,
+$G_{3,1}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\qquad G_{3,2}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix},\qquad G_{3,3}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix}$$\qquad$ $G_{3,4}=\frac{1}{\sqrt{6}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}$.
+For the first two POVMs ($\alpha=1,2$), we choose the mutually unbiased measurements Kalev
+$E_{\alpha,k}=\left\{\begin{aligned} &\frac{1}{3}\mathbb{I}_{3}+t_{\alpha}\left[G_{1,1}+G_{1,2}-\sqrt{3}(1+\sqrt{3})G_{\alpha,k}\right],\quad k=1,2,\\ &\frac{1}{3}\mathbb{I}_{3}+t_{\alpha}(1+\sqrt{3})(G_{1,1}+G_{1,2}),\qquad k=3.\end{aligned}\right$.
+(44)
+The last POVM is a $(1,5)$-POVM SIC-MUB with
+$E_{3,k}=\left\{\begin{aligned} &\frac{1}{5}\mathbb{I}_{3}+t_{3}\left[G_{3,1}+G_{3,2}+G_{3,3}+G_{3,4}-\sqrt{5}(1+\sqrt{5})G_{\alpha,k}\right],\quad k=1,2,3,4,\\ &\frac{1}{5}\mathbb{I}_{3}+t_{3}(1+\sqrt{5})(G_{3,1}+G_{3,2}+G_{3,3}+G_{3,4}),\qquad k=5,\end{aligned}\right$.
+(45)
+The parameters $t_{\alpha}$ are chosen to be non-optimal – that is, there exist greater values of $t_{\alpha}$ that guarantee $E_{\alpha,k}\geq 0$. To simplify further calculations, we take
+$t_{1}=t_{2}=\frac{1}{3(1+\sqrt{3})},\qquad t_{3}=\frac{1}{10(1+\sqrt{5})}.$ (46)
+The corresponding generalized equiangular measurement follows from
+$P_{\alpha,k}=\gamma_{\alpha}E_{\alpha,k},$ (47)
+where $\gamma_{1}=\gamma_{2}$ and $\gamma_{3}=1-2\gamma_{1}$ with
+$\gamma_{1}=\sqrt{5}-2.$ (48)
+By construction, this GEAM is a conical 2-design with $S=9/4-\sqrt{5}$.
+The most general $3\times 3$ rotation matrix that preserves the maximally mixed vector is the cyclic matrix MUBs
+$\mathcal{R}(\theta)=\begin{pmatrix}c_{1}(\theta)&c_{2}(\theta)&c_{3}(\theta)\\ c_{3}(\theta)&c_{1}(\theta)&c_{2}(\theta)\\ c_{2}(\theta)&c_{3}(\theta)&c_{1}(\theta)\end{pmatrix}$ (49)
+with
+$c_{1}(\theta)=\frac{1}{3}+\frac{2}{3}\cos(\theta),\qquad c_{2}(\theta)=\frac{1}{3}+\frac{2}{3}\cos(\theta-2\pi/3),\qquad c_{3}(\theta)=\frac{1}{3}+\frac{2}{3}\cos(\theta+2\pi/3).$ (50)
+In the following examples, we take
+$\mathcal{O}^{(1)}=\mathcal{R}(2\pi/3),\qquad\mathcal{O}^{(2)}=\mathcal{R}(-2\pi/3),$ (51)
+and the $5\times 5$ permutation matrix
+$\mathcal{O}^{(3)}=\begin{pmatrix}0&1&0&0&0\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&0&1&0&0\\ 0&0&0&0&1\end{pmatrix}$. (52)
+Examples of PPT states $\rho$ detected via a given witness $W$ are constructed by demanding that $\rho$ has the same form as $W$.
+Then, conditions $\rho\geq 0$ and $\rho^{\Gamma}\geq 0$ are applied, and the remaining free parameters are chosen so that $\mathrm{Tr}(\rho W)<0$.
+Example 1.
+For $K=N=3$ and $L=2$, we construct the entanglement witness
+$W=(2\widetilde{C}_{2}-\widetilde{C}_{3})\mathbb{I}_{3}\otimes\mathbb{I}_{3}+J_{1}-J_{2}-J_{3},$ (53)
+where
+$\widetilde{C}_{2}=\frac{2}{3}S+\frac{\gamma_{2}^{2}}{3}+\frac{\gamma_{3}^{2}}{5}.$ (54)
+The explicit form of $W$ reads
+$W=\left[\begin{array}[]{c c c|c c c|c c c}x&\cdot&\cdot&\cdot&iu&\cdot&\cdot&\cdot&\cdot\\ \cdot&y&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&z&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&y&\cdot&\cdot&\cdot&\cdot&\cdot\\ -iu&\cdot&\cdot&\cdot&z&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&x&\cdot&-iv&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&z&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&iv&\cdot&x&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&y\end{array}\right]$ (55)
+where
+$\begin{split}&x=\frac{(2-\sqrt{3})(7-3\sqrt{5})}{12(3+\sqrt{5})},\qquad y=\frac{1}{6}(9-4\sqrt{5}),\qquad z=\frac{(2+\sqrt{3})(7-3\sqrt{5})}{12(3+\sqrt{5})},\\ u&=\frac{\sqrt{3}}{4}(9-4\sqrt{5}),\qquad v=\frac{1}{4\sqrt{161+72\sqrt{5}}},\end{split}$ (56)
+and the dots represent zeros in the matrix entries. This witness is indecomposable as it detects the rank-8 PPT entangled state
+$\rho=\frac{1}{300}\left[\begin{array}[]{c c c|c c c|c c c}43&\cdot&\cdot&\cdot&-2i\sqrt{258}&\cdot&\cdot&\cdot&\cdot\\ \cdot&33&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&24&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&33&\cdot&\cdot&\cdot&\cdot&\cdot\\ 2i\sqrt{258}&\cdot&\cdot&\cdot&24&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&43&\cdot&6i\sqrt{22}&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&24&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&-6i\sqrt{22}&\cdot&43&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&33\end{array}\right],$ (57)
+which can be shown by calculating the expectation value
+$\mathrm{Tr}(\rho W)\simeq-2.22\cdot 10^{-5}.$ (58)
+Example 2.
+This time, we fix $K=N=3$ and $L=1$. The entanglement witness is constructed as follows,
+$W=(2\widetilde{C}_{1}-\widetilde{C}_{3})\mathbb{I}_{3}\otimes\mathbb{I}_{3}+J_{1}-J_{2}+J_{3},$ (59)
+where
+$\widetilde{C}_{1}=\frac{2}{3}S+\frac{\gamma_{2}^{2}}{3}.$ (60)
+We find that the matrix representation of this witness is of the form
+$W=\left[\begin{array}[]{c c c|c c c|c c c}z&\cdot&\cdot&\cdot&iu&\cdot&\cdot&\cdot&\cdot\\ \cdot&y&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&x&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&y&\cdot&\cdot&\cdot&\cdot&\cdot\\ -iu&\cdot&\cdot&\cdot&x&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&z&\cdot&-iv&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&x&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&iv&\cdot&z&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&y\end{array}\right]$ (61)
+with the parameters given by eq. (56). Again, $W$ is indecomposable, allowing us to detect the rank-7 PPT entangled state
+$\rho=\frac{1}{750}\left[\begin{array}[]{c c c|c c c|c c c}31&\cdot&\cdot&\cdot&-5i\sqrt{186}&\cdot&\cdot&\cdot&\cdot\\ \cdot&69&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&150&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&69&\cdot&\cdot&\cdot&\cdot&\cdot\\ 5i\sqrt{186}&\cdot&\cdot&\cdot&150&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&31&\cdot&-31i&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&150&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&31i&\cdot&31&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&69\end{array}\right].$ (62)
+The expectation value of the witness $W$ in this state is equal to
+$\mathrm{Tr}(\rho W)\simeq-8.05\cdot 10^{-5}.$ (63)
 5 Separability criteria
@@
 $\mathcal{P}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big[\rho(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big].$ (64)
-The separability conditions follow in terms of its trace or trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$, and as it turns out they depend on the maximal indices of coincidence $\widetilde{\mathcal{C}}^{A}$, $\widetilde{\mathcal{C}}^{B}$ associated with $\{P_{\alpha,k}^{A}\}$ and $\{P_{\alpha,k}^{B}\}$, respectively.
+The first separability condition follows in terms of its trace, and as it turns out it depends on the maximal indices of coincidence $\widetilde{\mathcal{C}}^{A}$, $\widetilde{\mathcal{C}}^{B}$ associated with $\{P_{\alpha,k}^{A}\}$ and $\{P_{\alpha,k}^{B}\}$, respectively.
 Proposition 5.
@@
 From the linearity of the trace, this result extends from product states to all separable states. ∎
+Let us check the detection properties of this separability condition on the example of the Werner states Werner
+$\rho(\phi)=\frac{1}{d(d^{2}-1)}\Big[(d-\phi)\mathbb{I}_{d^{2}}+(d\phi-1)\mathbb{F}_{d}\Big],$ (67)
+where $1\leq\phi\leq 1$ and $\mathbb{F}_{d}$ is the flip operator.
+These states are entangled for $\phi<0$ and separable for $\phi\geq 0$.
+Assuming that the same GEAM $P_{\alpha,k}$ is used on both subsystems, one gets
+$\begin{split}\mathrm{Tr}\mathcal{P}&=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}\mathcal{P}_{\alpha,k;\alpha,k}=\sum_{\alpha=1}^{N}\frac{a_{\alpha}^{2}M_{\alpha}}{d(d^{2}-1)}[d-\phi+(d\phi-1)b_{\alpha}]\\ &=\frac{1}{d(d^{2}-1)}\left[d^{2}\mu(d-\phi)+(d\phi-1)\sum_{\alpha=1}^{N}a_{\alpha}^{2}b_{\alpha}M_{\alpha}\right].\end{split}$ (68)
+Now, remembering that $P_{\alpha,k}$ form a conical 2-design, we find that
+$S=a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=a_{\alpha}^{2}\frac{M_{\alpha}(db_{\alpha}-1)}{d(M_{\alpha}-1)}.$ (69)
+After multiplying both sides by $d(M_{\alpha}-1)$ and taking a sum over all $\alpha$, one arrives at
+$dS\sum_{\alpha=1}^{N}(M_{\alpha}-1)=\sum_{\alpha=1}^{N}a_{\alpha}^{2}M_{\alpha}(db_{\alpha}-1),$ (70)
+which further reduces to
+$d(d^{2}-1)S=d\sum_{\alpha=1}^{N}a_{\alpha}^{2}b_{\alpha}M_{\alpha}-d^{2}\mu.$ (71)
+Using the above equality in eq. (68) results in
+$\mathrm{Tr}\mathcal{P}=\mu+\frac{S}{d}(d\phi-1).$ (72)
+Hence, the separability criterion from Proposition 5 produces
+$\mu+\frac{S}{d}(d\phi-1)\leq\frac{d-1}{d}S+\mu,$ (73)
+which is equivalent to $\phi\leq 1$. This means that any conical 2-design GEAM detects all separable Werner states but no entanglement.
+Now, we take the isotropic states iso
+$\rho_{\rm iso}(\theta)=\frac{1-\theta}{d^{2}-1}(\mathbb{I}_{d^{2}}-P_{+})+\theta P_{+},$ (74)
+where $0\leq\theta\leq 1$, and $P_{+}=(1/d)\sum_{m,n=1}^{d}|m\rangle\langle n|\otimes|m\rangle\langle n|$ is a maximally entangled state.
+It is known that $\rho_{\rm iso}(\theta)$ are entangled for $\theta>1/d$ and separable for $\theta\leq 1/d$.
+Taking $P_{\alpha,k}^{A}=P_{\alpha,k}$ and $P_{\alpha,k}^{B}=\overline{P}_{\alpha,k}$ results in
+$\mathrm{Tr}\mathcal{P}=\sum_{\alpha=1}^{N}\frac{a_{\alpha}^{2}M_{\alpha}}{d(d^{2}-1)}\Big[d(1-\theta)+(d^{2}\theta-1)b_{\alpha}\Big]=\mu+\frac{S}{d}(d^{2}\theta-1),$ (75)
+where we used eq. (71) to simplify the formula. According to Proposition 5, $\rho_{\rm iso}(\theta)$ is separable if
+$\mu+\frac{S}{d}(d^{2}\theta-1)\leq\frac{d-1}{d}S+\mu,$ (76)
+or $\theta\leq 1/d$. If $\theta>1/d$, this state is entangled. Hence, any conical 2-design GEAM detects all separable and entangled isotropic states.
+Alternative separability conditions involve the trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$.
 Proposition 6.
@@
 If a bipartite state $\rho$ is separable, then
-$\|\mathcal{R}\|_{{\rm{tr}}}=\sqrt{\widetilde{\mathcal{C}}^{A}-\mathcal{C}^{A}(\rho^{A})}\sqrt{\widetilde{\mathcal{C}}^{B}-\mathcal{C}^{B}(\rho^{B})},$ (49)
+$\|\mathcal{R}\|_{{\rm{tr}}}\leq\sqrt{\widetilde{\mathcal{C}}^{A}-\mathcal{C}^{A}(\rho^{A})}\sqrt{\widetilde{\mathcal{C}}^{B}-\mathcal{C}^{B}(\rho^{B})},$ (79)
 where $\mathcal{C}^{A}(\rho^{A})$ and $\mathcal{C}^{B}(\rho^{B})$ are indices of coincidence for the marginal states of $\rho$.
@@
 Moreover, our results allow to simplify these criteria significantly by proving that they depend only on the index of coincidence for the marginal state and its upper bound.
+Example 3.
+Consider a bipartite qutrit state $\rho=p\rho_{x}+(1-p)P_{+}$ that is a mixture of Horodecki’s bound entangled state HORODECKI1997333
+$\rho_{x}=\frac{1}{1+8x}\left[\begin{array}[]{c c c|c c c|c c c}x&\cdot&\cdot&\cdot&x&\cdot&\cdot&\cdot&x\\ \cdot&x&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&x&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&x&\cdot&\cdot&\cdot&\cdot&\cdot\\ x&\cdot&\cdot&\cdot&x&\cdot&\cdot&\cdot&x\\ \cdot&\cdot&\cdot&\cdot&\cdot&x&\cdot&\cdot&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\frac{1+x}{2}&\cdot&\frac{\sqrt{1-x^{2}}}{2}\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&x&\cdot\\ x&\cdot&\cdot&\cdot&x&\cdot&\frac{\sqrt{1-x^{2}}}{2}&\cdot&\frac{1+x}{2}\end{array}\right],$ (90)
+$0<x<1$, and the maximally entangled state $P_{+}=(1/3)\sum_{m,n=1}^{3}|m\rangle\langle n|\otimes|m\rangle\langle n|$.
+Then, for this quantum state, construct the correlation matrix $\mathcal{P}$ with $P_{\alpha,k}^{A}=P_{\alpha,k}$ and $P_{\alpha,k}^{B}=\overline{P}_{\alpha,k}$, using the GEAM $\{P_{\alpha,k}\}$ given by eqs.
+(4448).
+Numerical calculations show the range of parameter $p$ for which entanglement of $\rho$ is detected for a fixed $x$ – that is, when the sufficient separability condition is broken.
+The results are presented in Table 1.
+Note that $\widetilde{C}$ is the upper bound for the index of coincidence corresponding to $P_{\alpha,k}$, and
+$\widetilde{C}_{AB}=\sqrt{\widetilde{\mathcal{C}}-\mathcal{C}(\rho^{A})}\sqrt{\widetilde{\mathcal{C}}-\mathcal{C}(\rho^{B})}.$ (91)
+Our results indicate that Proposition 7 detects the most and Proposition 5 the least entanglement.
+The differences in the range of $p$ decrease with the increase of the parameter $x$.
+Comparing to analogical calculations in ref.
+Shen , it is straightforward to see that our choice of a GEAM is able to detect more entanglement than MUMs and general SIC POVMs.
+Value of $x$ Entanglement criterion Range of $p$
+$x=0.1250$ $\mathrm{Tr}\mathcal{P}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9014$
+$\|\mathcal{P}\|_{\rm tr}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9310$
+$\|\mathcal{R}\|_{\rm tr}>\widetilde{\mathcal{C}}_{AB}$ $0\leq p\leq 0.9605$
+$x=0.3750$ $\mathrm{Tr}\mathcal{P}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9624$
+$\|\mathcal{P}\|_{\rm tr}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9722$
+$\|\mathcal{R}\|_{\rm tr}>\widetilde{\mathcal{C}}_{AB}$ $0\leq p\leq 0.9791$
+$x=0.6250$ $\mathrm{Tr}\mathcal{P}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9846$
+$\|\mathcal{P}\|_{\rm tr}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9884$
+$\|\mathcal{R}\|_{\rm tr}>\widetilde{\mathcal{C}}_{AB}$ $0\leq p\leq 0.9906$
+$x=0.8750$ $\mathrm{Tr}\mathcal{P}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9961$
+$\|\mathcal{P}\|_{\rm tr}>\widetilde{\mathcal{C}}$ $0\leq p\leq 0.9970$
+$\|\mathcal{R}\|_{\rm tr}>\widetilde{\mathcal{C}}_{AB}$ $0\leq p\leq 0.9975$
+Table 1: Range of the parameter $p$ for which entanglement of $\rho$ is detected by the criteria from Propositions 5–7.
 6 Conclusions
@@
 Interestingly, these witnesses are parametrized solely by two partial indices of coincidence.
+Examples for bound entanglement detection had been provided.
 Next, we used the generalized equiangular measurements to construct separability criteria for bipartite states with subsystems of arbitrary dimensions.
@@
 Construction of witnesses and separability criteria from measurements that are not conical 2-designs remains as an open question.
+Another possible direction is to generalize the presented formalism to multipartite systems and subsystems of unequal dimensionality.
 7 Acknowledgements
@@
 For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
+8 Author Contribution
+K.S. is solely responsible for the entire research process.